OpenStudy (anonymous):

HYPOTHESIS Calculate the test statistic. Using α=0.05, determine the critical region and hence write the appropriate decision rule. Table 1: Average fees for a Full-Service Funeral (n=36) 7.4 9.4 5.3 8.4 7.5 6.5 6.2 8.3 6.7 11.6 6.3 5.9 6.7 5.8 5.2 6.4 6 7.4 7.2 6.6 6.3 5.3 6.6 5.6 8.4 7.2 7.4 5.8 6.3 6.1 7 7.2 6.1 5.4 7.4 6.6

OpenStudy (anonymous):

Can you calculate the mean? YES Can you calculate the Standard Deviation? YES Can you calculate the Z score? YES (look it up in a Normal Distribution Table for alpha=0.05 two tails or 0.025 two tails). (I would use Student-t distribution thoug with 35 degrees of freedom because we do not know the true standard deviation of the population and need to to use the sample's) The interval is: \[[mean-Z_{\alpha/2}*\sigma/\sqrt{n},mean+Z_{\alpha/2}*\sigma/\sqrt{n}]\]

OpenStudy (anonymous):

Z has to be calculated for 0.05 (two tails) or 0.025 (one tail) It was wrong in my previous post